the wavelength of a spectral line for an electronic transition is inversely related to
a. the number of electrons undergoing the transition
b. the nuclear charge of the atom
c. the difference in the energy of the energy levels involved in the transition
d. the velocity of the electron undergoing the transition

Grade:upto college level

2 Answers

Navjyot Kalra
9 years ago
Transition energy (∆E) = kZ2 (1/(n_1^2 ))- (1/(n_2^2 )) = hc/λ Ie, ∆E ∝ 1/λ
Gazal chanda
6 years ago
Acc to rydberg formula we can say tht wave no is inverse of wavelengthAnd waveno is directly proportional to z which is atomic no or can say no of protondSo wavelength of a spectral line is inversely proportional to nuclear charge of atom( no of protond or atomic no) So b part is ansC part is also the correct answer bcoz acc to rydberg formula 1/ wavelength when goes to right side will came up and diff in energy levels came downside on rhs.... so they are inversely proportional .alsoHence ans of ques is both b as well as c part